OpenStudy (marigirl):

Integration question:

OpenStudy (marigirl):

\[-cosec(6x)\cot(6x)\]

OpenStudy (marigirl):

\[-cosec(6x)\cot(6x)\] \[\frac{ -1 }{ \sin(6x) }*\frac{ \cos (6x) }{\sin (6x)}\] Integrate \[\frac{ -\cos(6x) }{ \sin^2(6x) }\]

OpenStudy (misty1212):

HI!!

OpenStudy (marigirl):

Hi Misty1212 I need guidance from here plzz

OpenStudy (misty1212):

do you know a function whose derivative is \[\csc(x)\cot(x)\]?

OpenStudy (misty1212):

hint, you should !

OpenStudy (misty1212):

if you do not, i can tell you, but i assure you it is in your book under the section of "derivatives of trig functions"

OpenStudy (marigirl):

if y=cosec(x) then y'=-cosec(x)cot(x)

OpenStudy (misty1212):

ok you got it

OpenStudy (misty1212):

of course you have to adjust for the \(6x\) but that is a mental u - sub divide by 6 to take care of that

OpenStudy (misty1212):

it it clear what i mean?

OpenStudy (misty1212):

you need to change the sign, and divide by 6 to get the derivative you want, i.e. \[-\frac{1}{6}\csc(6x)\] will work

OpenStudy (marigirl):

yes because when you integrate you tend to divide by the power multiplied by the coefficient of x?

OpenStudy (misty1212):

yes it is the chain rule backwards

OpenStudy (misty1212):

the check is easy right? take the derivative of \[-\frac{1}{6}\csc(6x)\] and see that you get the integrand \[\csc(6x)\cot(66x)\]

OpenStudy (misty1212):

oops too many sixes, but you get the idea

OpenStudy (marigirl):

this is a once step thing which is fantastic thank you! but if for some reason I decided to make things complicated and go \[\frac{ -\cos(6x) }{ \sin^2(x) }\] can i use \[\int\limits \frac{ f'(x) }{ f(x) } \] rule somehow?

OpenStudy (misty1212):

you still need the mental u - sub

OpenStudy (misty1212):

and it would not be what you wrote, because of the square in the denominator

OpenStudy (misty1212):

it would be more like \[\int\frac{f'(x)}{f^2(x)}dx\]

OpenStudy (marigirl):

sorry i meant \[\frac{ -\cos(6x) }{ \sin^2(6x) }\]

OpenStudy (marigirl):

Sorry I didnt think I made myself very clear Could I possibly use \[A \int\limits \frac{ f'(x) }{ f(x) } dx= A \ln \left| fx \right| +c\]

OpenStudy (misty1212):

no, as i said, you are missing the square

OpenStudy (misty1212):

\[\int\frac{f'(x)}{f^2(x)}dx\]j you will not get the log

OpenStudy (marigirl):

oh yea i get it now .. best method is to use the basic rules outlined . Thanks Heaps! I cant believe i made it so complicated lol

OpenStudy (misty1212):

lol it is easy to do with integration look for the simple first \[\color\magenta\heartsuit\]